OpenStudy (anonymous):

can anyone help me with this? Heres the question: The student council is selling cases of gift cards as a fundraiser. The revenue, r(x), in dollars, can be modelled by the function R(x)= -25x(squared) +100x +1500, where x is the number of cases of gift cards sold. How many cases must be sold to maximize the revenue?

OpenStudy (amistre64):

i can help...maybe

OpenStudy (amistre64):

you doing this for calculus right?

OpenStudy (anonymous):

mixed math grade 11

OpenStudy (amistre64):

do you know what a "derivative" is yet?

OpenStudy (anonymous):

not yet, no

OpenStudy (amistre64):

Its alot quicker with the "derivative"; but we can do it without... you up for it?

OpenStudy (anonymous):

sure *!

OpenStudy (amistre64):

have you gone over the quadratic equation or "completeing the square" yet?

OpenStudy (anonymous):

yes, but i do not remember how to complete the square unfortunately

OpenStudy (amistre64):

thats ok, completing the square is PROOF that the quadratic formula works...

OpenStudy (anonymous):

alright

OpenStudy (amistre64):

the only thing we are concerned about with the quadratic formula is what we get for -b/(2a) cause that gives us the x value for where this equation bends over itself. It looks like aupside down bowl. Ok?

OpenStudy (anonymous):

okay i understand that

OpenStudy (amistre64):

a = first = -25; b = middle = 100 Do you agree with this?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

-b/2a: -(100)/(2(-25)) = -50/-25 = -2/-1 = 2. Am I right?

OpenStudy (anonymous):

wait, where did u get the 2?

OpenStudy (amistre64):

Im old now, so I could be forgetting something :)

OpenStudy (amistre64):

which one? the -b/2a..that 2?

OpenStudy (anonymous):

how does the quadratic formula go again?

OpenStudy (amistre64):

Its got two parts to it; or rather, we can seperate it into 2 parts like this: -b/2a +- sqrt(b^2-4ac)/2a

OpenStudy (anonymous):

what is the slash?

OpenStudy (amistre64):

1/2 means 1 divided by 2 .... which means one-half..... but for now, just see it as a fraction bar

OpenStudy (anonymous):

alright, i think i understand now. so just substitute the numbers into the quadratic formula and i shall get 2? that is the answer in the back of the book

OpenStudy (amistre64):

-b (+-) sqrt(b^2-4ac) --- ------------ 2a 2a Does this help?

OpenStudy (amistre64):

Yes, but the only thing we care about for finding the maximum amount is what happens at x = -b/2a

OpenStudy (anonymous):

so basically just use x= -b divided by 2a and i will get the answer of 2?

OpenStudy (amistre64):

Lets find out, cause I could have messed it up in my head :) -100 divided by 2(-25) = what? -100 divided by -50 = 2 right?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

Now that we now the value of "x" that will give us the most stuff, lets plug it into the original equation to figure out HOW MUCH stuff we get..... you agree?

OpenStudy (anonymous):

yes i think so

OpenStudy (amistre64):

R(x)= -25(2)^2 +100(2) +1500 R(x)= -25(4) + 200 + 1500 R(x)= -100 + 200 + 1500 R(x) = 100 +1500 R(x) = 1600 Do you agree?

OpenStudy (anonymous):

yes i agree

OpenStudy (amistre64):

the derivative of: R(x)= -25x^2 +100x +1500 is -50x +100. And when the derivative = 0 we get the max amount: when x=100/50 = 2 the derivative = 0. Do you see that?

OpenStudy (anonymous):

thats kind of confusing... but you helped me find out how to get the right answer it really helped thanks!

OpenStudy (amistre64):

lol..... youre welcome :)